• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name L. Zhu

    Log entry time 10:23:01 on April25,2007

    Entry number 199988

    keyword=controlled access

    Controlled access was asked to reset left Q3. Bob is working on the missing TDC module.

    For the moment, The helium flow to our magnets is still not stable due to the work on hall C HMS. We lost our magnets twice. Cryo group is turning on another compressor for Hall C. Then we will reset the mangets again.